OpenStudy (anonymous):

evaluate the following: log5 (1/125)

myininaya (myininaya):

[ln1/125]/[ln5]=[ln125^(-1)]/[ln5]=[ln5^(-3)]/[ln5] see if that helps

myininaya (myininaya):

what do we do next

myininaya (myininaya):

you can bring that -3 down right? does anything cancel?

OpenStudy (anonymous):

get rid of the exponent?

OpenStudy (anonymous):

oh yes i was rite. um the 5's cancel rite?

myininaya (myininaya):

so we have -3(ln5)/(ln5)=-3 the ln5 cancels

OpenStudy (anonymous):

so where left with -3?

OpenStudy (anonymous):

hello?

myininaya (myininaya):

?

OpenStudy (laloo):

We can write it as \[\log_{5}1-\log_{5}125 \] .................. (1) then we know that logarithm of 1 with any base is zero so the (1) becomes \[0-\log_{5}125\] ................... (2) now we can write this as \[-\log_{5}5^{3}\] ................... (3) now this becomes \[-3\times \log_{5}5 \] ................... (4) and as we know \[\log_{5}5=1 \] ...................... (5) so answer is -3

OpenStudy (anonymous):

i got the rite answer :) thankyou!

myininaya (myininaya):

very good thats what we got above

OpenStudy (laloo):

But in more simple way. :)